OpenStudy (anonymous):

Jeremy Wants To Buy A New Computer. The Saleswoman Says That He Can Make A Down Payment And Then Pay For The Computer In Installments. Here's A Formula That Describes This Scenario: x=t-yz Where X=Amount Down Y=Money Each Month Z=Number Of Months T=Total Price Rewrite The Formula To Solve For The Amount Of Money Jeremy Must Pay Each Month.

OpenStudy (anonymous):

@ajprincess PLEASE HELP

OpenStudy (ajprincess):

Amount of money that Jeremy must pay each month is y. So solve ur equation for y. can u do it @DamnImBad?

OpenStudy (anonymous):

Umm Im Confused Just A Little

OpenStudy (ajprincess):

Follow the steps. 1.Add yz both sides. 2.Subtract X both sides. 3.Divide both sides by z. Can u do nw?

OpenStudy (anonymous):

1.xzy

OpenStudy (anonymous):

Hold On Im Tryna Remember

OpenStudy (anonymous):

So Will The First Be Like This:x=yz/t

OpenStudy (anonymous):

@ajprincess

OpenStudy (anonymous):



OpenStudy (ajprincess):

X=t-yz 1.Add yz both sides. X+yz=t-yz+yz X+yz=t 2.Subtract X both sides. X+yz-X=t-X yz=t-X

OpenStudy (anonymous):

So Then

OpenStudy (anonymous):